OpenStudy (anonymous):

Find the equation of the line of intersection of the planes with the equations -3x+4y+z=1 and 5x-3z=5

OpenStudy (anonymous):

Please help me work it out step by step :(

OpenStudy (anonymous):

alright so you will have to use the cross product of the two normal vectors

OpenStudy (anonymous):

I did and I got -12i-4j-20k

OpenStudy (anonymous):

hm...how'd u get yours?

OpenStudy (anonymous):

wait...i got it

OpenStudy (amistre64):

practice makes perfect, good job so far cherio ;)

OpenStudy (anonymous):

haha..gotta love calc =D..and ya i got -12i-4j-20k after i set it up right

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

so this is your direction vector i believe

OpenStudy (anonymous):

now you need to to find a point that lie on this line

OpenStudy (anonymous):

you can do this by taking your two planes and setting x,y, or z to zero and finding a point

OpenStudy (anonymous):

okay so if x=0 I get y=2/3 and z=-5/3 , right?

OpenStudy (anonymous):

yes, so for the next part you have to make a vector with that point (this will be dotted with the direction vector from before)

OpenStudy (anonymous):

so typically for a line you just select the (x,y,z) to be your point (if anything doesn't make sense let me know) so you get (x-0, y-(2/3), z+(5/3)) as that vector

OpenStudy (anonymous):

now i dot it with any scalar of the direction vector and get L=(x, y-(2/3), z+(5/3))* t(-12,-4,-20)

OpenStudy (anonymous):

Any reason why you do (x-0,y-(2/3),z+(5/3)? and you can't simplify that any more so that would be the answer?

OpenStudy (anonymous):

well if you want to write in parametric or (something that starts with a s...symmetric?) you can simplify it a bit more..but as far at why x,y,z...um..i'll try to draw a picture.

OpenStudy (amistre64):

in case i missed it, since the second plane has no y value, wouldnt it be simper to make y=0 in the first to solve for x,z ?

OpenStudy (anonymous):